A 500.0 ml. Glass filled with air is placed into water up-side-down while
at 7.0°C. The water is heated to 77°C. How much air bubbles out from
under the glass?
1
Expert's answer
2020-12-14T14:36:50-0500

Find the volume of expanded air assuming the pressure is constant:


"\frac{V_1}{T_1}=\frac{V_2}{T_2},\\\space\\
V_2=V_1\frac{T_2}{T_1}=500\frac{77+273}{7+273}=625\text{ ml}."

The volume of the glass is still 500 ml, so, the volume of bubbles is 125 ml.